LNO section meeting on commissioning review and MD preparation

Europe/Zurich
Zoom Meeting ID
67926097998
Host
Rogelio Tomas Garcia
Useful links
Join via phone
Zoom URL

Meeting on commissioning review and MD preparation

News / Introduction
Rogelio reminds about internal IPAC schedules (sent to GL by today) and welcomes newcomers in this section. Elene Kravishvili works as a shorter term intern with Felix Charlier on the implementation of machine learning techniques for linear triplet optics corrections in the LHC. Satya Jagabathuni moved earlier this year to this section and Kyriacos Skoufaris joined as member for both LNO and NDC sections working on FCC and LHC. There is good progress in LHC commissioning as also discussed in the presentations today,
Laurent announces that this years International Computational Accelerator Physics Conference ICAP’24 scheduled for October is now open for registration.

LHC commissioning OMC,  overview 1,  Ewen, slides
Ewen presents an overview over the recent LHC optics measurement and correction activities. So far the have been 14 measurement shifts. They started at injection energy on March 8 until and continued until  currently April 9. They took place mostly outside working hours, in many cases during nights, covering both core activities as well as some additional (filler) activities when no other teams were available. The shifts were also very useful to check and in some cases adapt the OMC software tools (IP1 polarity change). Ewen shows a  list of the activities in time order and discusses and illustrates some of the highlights. Linear optics measurements at injection showed that the virgin beta beat has increased compared to 2023. Global corrections from last year were used at 2m and 1.2 m at top energy and found to work well. Four iterations (included energy corrections) were used to correct and incorporate trims in the ramp/squeeze. The new non-linear B4 correction compensates well the quadratic feed-down observed at IP1 in a crossing scan at beta* = 30 cm and improves the tune signal.
LHC commissioning OMC,  overview 2, Tobias, slides
Tobias continues the overview and also comments on what remains to be done. Early reports of an apparent ATLAS/CMS appear to in part originated in an ATLAS pile-up dependence that is more significant with fewer bunches. In a recent 1200 bunch fill, the ATLAS and CMS luminosities agreed very well. He shows on/off momentum measurements of the beta beat at the lowest beta* of 22 cm reached so far with flat optics in IP1/5, and comments that he hopes this can be used in operation after the technical stop. Measurements have also been performed to check that the wire compensator can be activated. A first set of optics measurements has also been done over easter for the ion optics, initially without crossing angles. One to two shifts will be needed to repeat them with crossing angles. Measurements with K-modulation have been done and will be analyzed. More work is planned to use the DOROS beam pickups and use them also in turn by turn measurements.

Experience with the anomalous Dy in Beam 2, Yannis Angelis,  slides
A series of measurements has been performed during the commissioning with the aim to better understand an anomalous vertical dispersions visible on beam2. The origin of this has not yet been identified. There are ideas that it may be related to powering errors or misalignment of MCS (spool piece correctors).  Yannis shows orbits and dispersions measured at injection. The measurements were repeated using different MCS, MQS and MS settings. Some differences were seen. Rogelio and Riccardo suggest to try to modify the plotting to make changes more visible. Rogelio comments that the proposed 60 degree lattice MD could also be of interest to check dispersion with different main quadrupole settings. 

Status of the LHC optics for the HL-LHC optics MD, Riccardo
Riccardo says he works with Björn Lindstrom and colleagues on the specification of the HL-LHC optics for the upcoming MD1, 2 sessions, planning to finalize this by the end of next week. Riccardo made a full implementation of his optics factory now all in Python and taking advantage of Xsuite for the matching, and applying this for the first time for actual commissioning. The flexibility of this method and the speed achieved in complex matching are very promising. Basic collimation settings have been defined. A TCDQ gap change is proposed but would need an access currently not foreseen.  Björn has ideas to modify IR3, 7 optics to further improve cleaning efficiency,  but would need more time (months) to work out all details. Riccardo asks when the beta* at flat top ( 2.5 m or more ? ) and end of squeeze ( ~ 37 cm )  need to be fixed to be ready for MD1,2. This should be discussed with OP (Matteo et al.), in a small meeting to be organized soon, like beginning of next week.

There are minutes attached to this event. Show them.